OpenStudy (anonymous):

Given the equation y − 3 = 1/2(x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. (A:x − 2y = −12) (B:y =1/2x) (C: y = 1/2x + 6) (D: y =1/2x + 9)

OpenStudy (anonymous):

Please help if anyone can :(

OpenStudy (anonymous):

well here we go y − 3 = 1/2(x + 6) ok lets remove parentheses and to do so we must multiply... y-3= 1/2x +one half of 6 that would be 3 so now we have... y − 3 = 1/2x + 3 now add 3 to both sides and you get... y = 1/2x + 6 Letter C

OpenStudy (anonymous):

Thank you sooooooo much!